A display for rolls of tape indicates that each roll contains 150 yards of tape. If the actual length of tape can vary by 2.5% of that amount, what is the range for the length of tape on a roll?
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52786)

.

From minimum value of  150 - 0.025*150 = 146.25 yards

 to  maximum value of  150 + 0.025*150 = 153.75 yards.

Solved.



Answer by math_tutor2020(3817)


Min = 150*(1-0.025) = 146.25 yards
Max = 150*(1+0.025) = 153.75 yards